### v1

# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

Abstract
Abstract
The launch of Grokipedia, an AI-generated encyclopedia developed by Elon Musk’s xAI, was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries via the large language model Grok. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study undertakes a large-scale computational comparison of 382 matched article pairs between Grokipedia and Wikipedia. Using metrics across lexical richness, readability, structural organization, reference density, and semantic similarity, we assess how closely the two platforms align in form and substance. The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and more variable structural depth. These findings suggest that AI-generated encyclopedic content currently mirrors Wikipedia’s informational scope but diverges in editorial norms, favoring narrative expansion over citation-based verification. The implications highlight new tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
1 Introduction
Online encyclopedias have become infrastructural to public knowledge, shaping how people learn about politics, science, culture, and current events. Wikipedia, launched in 2001, is the paradigmatic example: a nonprofit, volunteer-edited project governed by policies like Neutral Point of View (NPOV) and verifiability. Despite its success, Wikipedia has long faced questions about bias and reliability. [^1] For a survey of evidence and debates on reliability and bias, see e.g., [4, 8, 17, 18].
In late 2025, xAI launched Grokipedia, an AI-generated encyclopedia positioned explicitly as an alternative to Wikipedia. [^2] See [20, 1, 19, 11, 13, 6, 10]. According to xAI’s framing and public statements, Grokipedia aims to “purge out the propaganda” and deliver “truthful” entries, with content generated and “fact-checked” by the Grok language model rather than a community of human editors.[20, 11, 10] At launch (October 27 2025), coverage reported approximately 800–900 k entries, limited public editability (suggest-edits rather than direct editing), and claims of openness alongside ambiguity about licensing and source code availability.[20, 19, 1] Early reception focused on two themes: (i) accusations that many Grokipedia pages were adapted or copied from Wikipedia (sometimes nearly verbatim), and (ii) concerns about ideological framing and AI-specific failure modes (e.g., hallucinations, dataset or alignment bias).[13, 6, 11, 10]
The launch has intensified an older debate: is Wikipedia biased, and if so, can an AI-first encyclopedia fare any better? Empirically, scholarly work has documented multiple kinds of bias in Wikipedia, including early left-leaning slant in U.S. politics entries that attenuated with maturation,[4] topical and coverage biases,[8] and sentiment or framing asymmetries in political pages.[17, 18] Meanwhile, contemporary studies show that large language models (LLMs) exhibit measurable political and social biases that vary by model and prompt context, raising separate concerns about any AI-generated encyclopedia.[22, 15, 3, 9]
Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We approach this by pairing articles across platforms and comparing them along four dimensions: (1) lexical similarity (e.g., TF–IDF cosine, $n$-gram overlap), (2) semantic similarity (embedding cosine, BERTScore), (3) structural similarity (headings and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, and part-of-speech distributions). Beyond similarity, we also examine descriptive characteristics—article length, readability, reference density, and link structure—since editorial workflows (human vs. AI-generated) plausibly shape article form as much as content.
To construct a meaningful comparison, we began from Wikipedia’s 416 most-edited articles, which prior research shows are strongly associated with high-controversy or high-salience topics.[25] Among these, 382 titles had valid, non-trivial matches on both platforms.
This comparison provides empirical evidence to evaluate current claims about the distinctiveness—or sameness—of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s content and structure, that would suggest an AI-mediated transformation with limited editorial divergence. Conversely, systematic differences in length, style, citation density, or framing could reveal an emerging AI editorial logic distinct from Wikipedia’s community governance and policy regime. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the future of reference knowledge online. [^3] For background critiques and expectations about Grokipedia’s neutrality and provenance, see [11, 6, 10, 23].
2 Methods
2.1 Sampling and Data Collection
Our analysis began with the selection of the 416 most-edited English-language Wikipedia articles, representing topics with the highest cumulative edit counts as of October 2025. Previous research has shown that highly edited articles are strongly correlated with controversy, salience, and social polarization [25]. This sampling strategy thus prioritizes articles that are both textually substantial and socioculturally significant.
For each title, we retrieved the corresponding entries from Wikipedia and GrokiPedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title>
and https://grokipedia.com/page/<Title>. HTML pages were downloaded via the requests library (Python 3.11) using polite delays and standard user-agent headers, and parsed with BeautifulSoup4. After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained for analysis. Out of 416 target titles, 382 matched article pairs satisfied these criteria and formed the final analytical sample.
2.2 Text Cleaning and Feature Extraction
We employed a host-aware parsing strategy designed to preserve substantive content while removing non-textual elements. For Wikipedia, extraction was restricted to the #mw-content-text container, excluding elements such as infoboxes, hatnotes, metadata, navboxes, and tables. For GrokiPedia, a less aggressive parser was used to retain the full model-generated narrative text, including embedded lists and inline references. Scripts, styles, and navigation elements were removed in both cases.
Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. We then computed a comprehensive set of descriptive, structural, and stylistic metrics:
  • Structural features: counts of paragraphs, headings (h1–h4), hyperlinks, images, and references; visible character counts; and derived ratios such as references, links, and headings per 1,000 words.
  • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
  • Lexical density and reading time: unique-to-total word ratio and estimated reading time assuming 200 words per minute.
All features were computed independently for both platforms and labeled with prefixes a_ (GrokiPedia) and b_ (Wikipedia).
2.3 Similarity Measures Between Platforms
To quantify the degree of alignment between matched GrokiPedia and Wikipedia articles, we computed a suite of nine similarity measures grouped into four conceptual domains:
  1. 1.
    Lexical similarity: cosine similarity of TF–IDF vectors (1–2 grams, English stop-words removed) and unigram Jaccard index.
  2. 2.
    N-gram overlap: overlap coefficients for 1-, 2-, and 3-gram sequences, capturing local phrase reuse.
  3. 3.
    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14].
  4. 4.
    BERTScore F1: contextual similarity of the first 50 sentences per article using
    bert-score [26].
  5. 5.
    Structural similarity: composite index combining (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths, weighted 0.6 and 0.4 respectively.
  6. 6.
    Stylistic similarity: normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition), transformed to a 0–1 similarity scale.
A composite similarity index was calculated as the mean of the lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
2.4 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. Descriptive statistics (mean $\pm$ SD) were calculated for all features. Differences between GrokiPedia and Wikipedia were tested using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were used to assess their interdependence and dimensional clustering. Visualizations—including histograms, boxplots, scatter plots, and correlation heatmaps—were generated with matplotlib and seaborn, then exported to PDF for direct inclusion in the LaTeX manuscript.
2.5 Limitations
The dataset covers only topics that are both heavily edited on Wikipedia and available on GrokiPedia. Because high edit frequency correlates with controversy and visibility [25], the sample likely overrepresents politically or culturally salient subjects. Moreover, GrokiPedia’s content-generation pipeline is opaque, potentially leading to stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose. Nevertheless, the scale and matched-pair design provide a robust basis for assessing how human and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
3 Results
3.1 Sample and Coverage
From the initial list of 416 highly edited Wikipedia titles, we successfully retrieved and matched 384 article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each). [^4] We use cleaned word counts extracted from the main content region, excluding navigation, infoboxes, and boilerplate.
3.2 Descriptive Differences Between Platforms
Figure 1 shows article length distributions (on a logarithmic scale). GrokiPedia entries are generally longer than their Wikipedia counterparts. Readability, measured with the Flesch–Kincaid grade, tends to be higher (i.e., more complex prose) on GrokiPedia (Figure 2). Table 1 summarises core descriptive measures, including clean word counts, readability, lexical diversity, and per-1,000-word densities of references, links, and headings.
Figure 1: Distribution of clean word counts by platform (log scale).
Figure 2: Readability (Flesch–Kincaid grade) by platform.
Table 1: Descriptive statistics (mean [SD]) across platforms (pairs with both articles and $\geq$500 clean words). $\Delta$ is GrokiPedia-Wikipedia. $p$ values from paired $t$-tests.
Measure GrokiPedia Wikipedia $\Delta$ $p$
Clean word count 65505 [74231] 18733 [22671] 46772 [62148] <0.001<0.001
Flesch–Kincaid grade 18.52 [5.81] 14.35 [4.90] 4.17 [5.34] <0.001<0.001
Lexical diversity 0.078 [0.037] 0.204 [0.045] 0.126-0.126 [0.056] <0.001<0.001
References per 1k words 0.45 [0.38] 1.26 [0.71] 0.81-0.81 [0.68] <0.001<0.001
Links per 1k words 3.18 [1.54] 4.01 [2.02] 0.83-0.83 [1.80] $0.012$
Headings (H2–H4) per 1k words 0.92 [0.41] 0.84 [0.39] 0.08 [0.37] $0.075$
3.3 Cross-Platform Similarity
We evaluated alignment between matched articles using lexical, semantic, structural, and stylistic metrics. Figure 3 presents the correlation matrix among these metrics. Lexical measures (TF–IDF, Jaccard, and $n$-gram overlap) tend to cluster with semantic similarity, while stylistic and structural measures form related but partially distinct dimensions. Table 2 reports the average values.
Figure 3: Correlation among similarity metrics (Pearson’s $r$).
Table 2: Average similarity metrics across matched pairs (mean [SD]).
Metric Mean [SD]
TF–IDF cosine 0.50 [0.22]
Jaccard (unigram) 0.32 [0.16]
Overlap (1-gram) 0.58 [0.23]
Overlap (2-gram) 0.33 [0.34]
Overlap (3-gram) 0.25 [0.37]
Semantic cosine 0.73 [0.12]
BERTScore (F1) 0.02 [0.24]
Structural similarity 0.31 [0.26]
Stylistic similarity 0.64 [0.08]
3.4 Distributions of Similarity Scores
To characterise the heterogeneity of alignment, Figure 4 shows the empirical distributions of all similarity scores across the matched pairs. Several metrics are unimodal and right-skewed (e.g., semantic cosine, BERTScore), suggesting generally high topical overlap. Others (e.g., higher-order $n$-gram overlap and structural similarity) exhibit broader or even multimodal shapes, consistent with clusters of near-verbatim matches versus paraphrastic or restructured articles.
Figure 4: Distributions of article-to-article similarity scores across metrics.
3.5 Does Length Drive Similarity?
Finally, we test whether longer Wikipedia entries are more semantically aligned with GrokiPedia. Figure 5 plots semantic cosine against the (log) Wikipedia word count. The association is weak-to-moderate and not monotonic, indicating that content focus and selection likely matter more than sheer length for semantic convergence.
Figure 5: Wikipedia length (log scale) vs. semantic similarity (cosine). A least-squares trend line is overlaid.
4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between GrokiPedia and Wikipedia across 382 matched article pairs. Despite their distinct production paradigms—community-driven versus AI-generated—the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.70, and stylistic similarity around 0.65, indicating that GrokiPedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. GrokiPedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels (greater syntactic complexity) but lower lexical diversity and reference density. This pattern suggests that GrokiPedia’s generation process elaborates on existing material—expanding text length and rhetorical flow—rather than producing substantively new or more rigorously sourced knowledge.
The similarity distributions reveal further nuance. While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are bimodal, suggesting two distinct regimes of generation. In some cases, GrokiPedia closely mirrors Wikipedia’s organization and register; in others, it diverges sharply, producing verbose or restructured narratives. This heterogeneity implies a hybrid authorship logic: part replication, part improvisation.
Elon Musk framed GrokiPedia as an antidote to Wikipedia’s “propaganda” and ideological bias [2, 7, 24]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many GrokiPedia articles are derived from Wikipedia—often copied or paraphrased—while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21]. Our quantitative results support this view: GrokiPedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density. In other words, GrokiPedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution. Rather than “purging bias,” the system appears to re-encode it—subtly and opaquely—within model parameters and prompt conditioning.
As argued by Yasseri [24], Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable. GrokiPedia inverts this model. Its authorship is singular, automated, and invisible—bias becomes latent, hidden within model weights and unseen editorial heuristics. This creates what might be termed an epistemic opacity paradox: GrokiPedia appears neutral because it lacks human editors, yet its underlying generative logic is uninspectable and unaccountable. The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 416 most-edited pages, of which 382 have counterparts on GrokiPedia. This selection likely overrepresents high-profile, contentious topics—those most prone to editing wars or ideological scrutiny [25]. Second, the similarity metrics employed—lexical, semantic, structural, and stylistic—quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while GrokiPedia’s generative parameters may change with future model retraining. Finally, GrokiPedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.
Despite its stated aim to “correct” Wikipedia’s ideological slant, GrokiPedia currently functions less as an epistemic alternative and more as a synthetic derivative of Wikipedia. It mirrors Wikipedia’s informational scope and linguistic tone but replaces community deliberation with algorithmic synthesis. The result is an encyclopedia that is fluent yet fragile—expansive in form but shallow in verifiability. In effect, GrokiPedia trades collective accountability for computational authority.
These findings extend beyond a single platform. As large language models increasingly mediate the production of knowledge, they risk conflating eloquence with accuracy and replacing transparency with automation. The challenge ahead is therefore not only technical—improving factuality or retrieval—but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge.
Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust. Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether GrokiPedia’s expansion of text corresponds to an expansion—or erosion—of understanding.
5 Acknowledgement
This publication has emanated from research supported in part by grants from Taighde Eireann – ´ Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.

### v2

@@
 Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear.
-This study undertakes a large-scale computational comparison of 382 matched article pairs between Grokipedia and Wikipedia.
+This study undertakes a large-scale computational comparison of 1,800 matched article pairs between Grokipedia and Wikipedia, drawn from the 2,000 most-edited Wikipedia pages.
 Using metrics across lexical richness, readability, structural organization, reference density, and semantic similarity, we assess how closely the two platforms align in form and substance.
-The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and more variable structural depth.
+The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and greater structural variability.
 These findings suggest that AI-generated encyclopedic content currently mirrors Wikipedia’s informational scope but diverges in editorial norms, favoring narrative expansion over citation-based verification.
@@
 Online encyclopedias have become infrastructural to public knowledge, shaping how people learn about politics, science, culture, and current events.
-Wikipedia, launched in 2001, is the paradigmatic example: a nonprofit, volunteer-edited project governed by policies like Neutral Point of View (NPOV) and verifiability.
-Despite its success, Wikipedia has long faced questions about bias and reliability.
+Wikipedia, launched in 2001, remains the paradigmatic example: a nonprofit, volunteer-edited project governed by policies such as Neutral Point of View (NPOV) and verifiability.
+Despite its success, Wikipedia has long faced questions about bias, reliability, and systemic underrepresentation.
 [^1] For a survey of evidence and debates on reliability and bias, see e.g., [4, 8, 17, 18].
-In late 2025, xAI launched Grokipedia, an AI-generated encyclopedia positioned explicitly as an alternative to Wikipedia.
+In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia.
 [^2] See [20, 1, 19, 11, 13, 6, 10].
-According to xAI’s framing and public statements, Grokipedia aims to “purge out the propaganda” and deliver “truthful” entries, with content generated and “fact-checked” by the Grok language model rather than a community of human editors.[20, 11, 10] At launch (October 27 2025), coverage reported approximately 800–900 k entries, limited public editability (suggest-edits rather than direct editing), and claims of openness alongside ambiguity about licensing and source code availability.[20, 19, 1] Early reception focused on two themes: (i) accusations that many Grokipedia pages were adapted or copied from Wikipedia (sometimes nearly verbatim), and (ii) concerns about ideological framing and AI-specific failure modes (e.g., hallucinations, dataset or alignment bias).[13, 6, 11, 10]
-The launch has intensified an older debate: is Wikipedia biased, and if so, can an AI-first encyclopedia fare any better?
-Empirically, scholarly work has documented multiple kinds of bias in Wikipedia, including early left-leaning slant in U.S.
-politics entries that attenuated with maturation,[4] topical and coverage biases,[8] and sentiment or framing asymmetries in political pages.[17, 18] Meanwhile, contemporary studies show that large language models (LLMs) exhibit measurable political and social biases that vary by model and prompt context, raising separate concerns about any AI-generated encyclopedia.[22, 15, 3, 9]
+According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[20, 11, 10] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800–900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[20, 19, 1] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[13, 6, 11, 10]
+The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better?
+Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[4] topical and coverage imbalances,[8] and framing asymmetries across ideology and gender.[17, 18] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[26] Parallel work on large language models (LLMs) shows measurable political and cultural biases that vary across model families and prompts,[22, 15, 3, 9] raising distinct questions about the epistemic reliability of AI-generated text.
 Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice?
-We approach this by pairing articles across platforms and comparing them along four dimensions: (1) lexical similarity (e.g., TF–IDF cosine, $n$-gram overlap), (2) semantic similarity (embedding cosine, BERTScore), (3) structural similarity (headings and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, and part-of-speech distributions).
-Beyond similarity, we also examine descriptive characteristics—article length, readability, reference density, and link structure—since editorial workflows (human vs.
-AI-generated) plausibly shape article form as much as content.
-To construct a meaningful comparison, we began from Wikipedia’s 416 most-edited articles, which prior research shows are strongly associated with high-controversy or high-salience topics.[25] Among these, 382 titles had valid, non-trivial matches on both platforms.
-This comparison provides empirical evidence to evaluate current claims about the distinctiveness—or sameness—of Grokipedia relative to Wikipedia.
-If Grokipedia largely reproduces Wikipedia’s content and structure, that would suggest an AI-mediated transformation with limited editorial divergence.
-Conversely, systematic differences in length, style, citation density, or framing could reveal an emerging AI editorial logic distinct from Wikipedia’s community governance and policy regime.
-By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the future of reference knowledge online.
-[^3] For background critiques and expectations about Grokipedia’s neutrality and provenance, see [11, 6, 10, 23].
+We address this by pairing articles across platforms and comparing them along four analytic dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), (3) structural similarity (section and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, part-of-speech composition).
+In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
+To construct a broad and representative corpus, we began with Wikipedia’s 2,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[26] Among these, approximately 1,800 titles had valid, nontrivial matches on Grokipedia.
+This enlarged sample enables more stable estimates of cross-platform differences than earlier pilot comparisons limited to a few hundred pairs.
+This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia.
+If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure.
+Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability.
+By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI.
+[^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [11, 6, 10, 23].
 2 Methods
 2.1 Sampling and Data Collection
-Our analysis began with the selection of the 416 most-edited English-language Wikipedia articles, representing topics with the highest cumulative edit counts as of October 2025.
-Previous research has shown that highly edited articles are strongly correlated with controversy, salience, and social polarization [25].
-This sampling strategy thus prioritizes articles that are both textually substantial and socioculturally significant.
+We analyzed the 2,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts.
+Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [26].
+This sampling strategy thus prioritizes articles that are both textually substantial and socially significant.
 For each title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title>
-and https://grokipedia.com/page/<Title>.
-HTML pages were downloaded via the requests library (Python 3.11) using polite delays and standard user-agent headers, and parsed with BeautifulSoup4.
-After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained for analysis.
-Out of 416 target titles, 382 matched article pairs satisfied these criteria and formed the final analytical sample.
+and
+https://Grokipedia.com/page/<Title>.
+HTML pages were downloaded via the requests library (Python 3.11) with polite delays and a standard user-agent header, then parsed using BeautifulSoup4.
+The extraction and analysis pipeline continuously wrote progress files (wiki_Grokipedia_similarity_live.csv) to enable incremental recovery and monitoring.
+After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained.
+Of the 2,000 target titles, 1,811 matched article pairs satisfied these criteria and formed the final analytical sample.
 2.2 Text Cleaning and Feature Extraction
-We employed a host-aware parsing strategy designed to preserve substantive content while removing non-textual elements.
-For Wikipedia, extraction was restricted to the #mw-content-text container, excluding elements such as infoboxes, hatnotes, metadata, navboxes, and tables.
-For GrokiPedia, a less aggressive parser was used to retain the full model-generated narrative text, including embedded lists and inline references.
-Scripts, styles, and navigation elements were removed in both cases.
-Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. We then computed a comprehensive set of descriptive, structural, and stylistic metrics:
+We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
+For Wikipedia, extraction was restricted to the #mw-content-text container and limited strictly to <p> and <li> elements.
+Elements such as infoboxes, hatnotes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
+For Grokipedia, a more adaptive extractor identified the main article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high link-density regions.
+Scripts, styles, and sidebars were removed in both cases.
+Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, structural, and stylistic metrics was computed:
   • Structural features: counts of paragraphs, headings (h1–h4), hyperlinks, images, and references; visible character counts; and derived ratios such as references, links, and headings per 1,000 words.
   • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
-  • Lexical density and reading time: unique-to-total word ratio and estimated reading time assuming 200 words per minute.
+  • Lexical density and reading time: ratio of unique to total words and estimated reading time assuming 200 words per minute.
 All features were computed independently for both platforms and labeled with prefixes a_ (Grokipedia) and b_ (Wikipedia).
 2.3 Similarity Measures Between Platforms
-To quantify the degree of alignment between matched GrokiPedia and Wikipedia articles, we computed a suite of nine similarity measures grouped into four conceptual domains:
+To quantify cross-platform alignment, we computed a suite of nine similarity measures grouped into four conceptual domains:
   1. 1.
@@
   3. 3.
-    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14].
+    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14], and contextual similarity via BERTScore F1 for the first 50 sentences per article [27].
   4. 4.
-    BERTScore F1: contextual similarity of the first 50 sentences per article using
-    bert-score [26].
-  5. 5.
-    Structural similarity: composite index combining (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths, weighted 0.6 and 0.4 respectively.
-  6. 6.
-    Stylistic similarity: normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition), transformed to a 0–1 similarity scale.
-A composite similarity index was calculated as the mean of the lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
+    Structural and stylistic similarity: a composite of (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths (weighted 0.6 and 0.4, respectively), together with normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
+A composite similarity index was calculated as the mean of lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
 2.4 Statistical and Comparative Analysis
@@
 Descriptive statistics (mean $\pm$ SD) were calculated for all features.
-Differences between GrokiPedia and Wikipedia were tested using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise.
-Spearman rank correlations among similarity metrics were used to assess their interdependence and dimensional clustering.
-Visualizations—including histograms, boxplots, scatter plots, and correlation heatmaps—were generated with matplotlib and seaborn, then exported to PDF for direct inclusion in the LaTeX manuscript.
+Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise.
+Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering.
+Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib and exported directly to PDF for inclusion in the manuscript.
 2.5 Limitations
 The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia.
-Because high edit frequency correlates with controversy and visibility [25], the sample likely overrepresents politically or culturally salient subjects.
-Moreover, GrokiPedia’s content-generation pipeline is opaque, potentially leading to stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose.
-Nevertheless, the scale and matched-pair design provide a robust basis for assessing how human and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
+Because edit frequency correlates with controversy and visibility [26], the sample likely overrepresents politically or culturally salient topics.
+In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose.
+Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
 3 Results
 3.1 Sample and Coverage
-From the initial list of 416 highly edited Wikipedia titles, we successfully retrieved and matched 384 article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each).
-[^4] We use cleaned word counts extracted from the main content region, excluding navigation, infoboxes, and boilerplate.
+From the expanded list of the 2,000 most-edited English Wikipedia titles, we successfully retrieved and matched 1,811 Grokipedia–Wikipedia article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each).
+[^4] Clean word counts were computed from visible prose only, excluding infoboxes, scripts, and navigation templates.
+This larger corpus substantially extends the coverage of the original 416-title sample while preserving topical breadth and diversity.
 3.2 Descriptive Differences Between Platforms
-Figure 1 shows article length distributions (on a logarithmic scale).
-GrokiPedia entries are generally longer than their Wikipedia counterparts.
-Readability, measured with the Flesch–Kincaid grade, tends to be higher (i.e., more complex prose) on GrokiPedia (Figure 2).
-Table 1 summarises core descriptive measures, including clean word counts, readability, lexical diversity, and per-1,000-word densities of references, links, and headings.
+Figure 1 compares article-length distributions on a logarithmic scale.
+Grokipedia entries remain systematically longer than their Wikipedia counterparts, though the gap narrows compared with the smaller sample.
+Readability, measured by the Flesch–Kincaid grade level (panel B of Figure 2), is again higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
+Table 1 summarises key descriptive statistics.
+Across the enlarged dataset, Grokipedia articles average about 14,200 words versus 9,400 on Wikipedia, and they exhibit lower lexical diversity and far fewer explicit references, links, and headings per thousand words.
+These contrasts reinforce that Grokipedia expands text length through elaboration rather than citation density or structural detail.
+Figure 2 visualises these platform differences (means $\pm$ SE) across all six measures reported in Table 1.
 Figure 1: Distribution of clean word counts by platform (log scale).
-Figure 2: Readability (Flesch–Kincaid grade) by platform.
-Table 1: Descriptive statistics (mean [SD]) across platforms (pairs with both articles and $\geq$500 clean words). $\Delta$ is GrokiPedia-Wikipedia. $p$ values from paired $t$-tests.
+Figure 2: Platform differences for key descriptives (means $\pm$ standard error).
+Panels show: (A) clean word count (log10), (B) Flesch–Kincaid grade, (C) lexical diversity (TTR), (D) references per 1k words, (E) links per 1k words, and (F) headings (H2–H4) per 1k words.
+Table 1: Descriptive statistics across matched article pairs (N=1,811N{=}1{,}811). Values are means [SD]; $p$ from paired $t$-tests.
 Measure Grokipedia Wikipedia $\Delta$ $p$
-Clean word count 65505 [74231] 18733 [22671] 46772 [62148] <0.001<0.001
-Flesch–Kincaid grade 18.52 [5.81] 14.35 [4.90] 4.17 [5.34] <0.001<0.001
-Lexical diversity 0.078 [0.037] 0.204 [0.045] 0.126-0.126 [0.056] <0.001<0.001
-References per 1k words 0.45 [0.38] 1.26 [0.71] 0.81-0.81 [0.68] <0.001<0.001
-Links per 1k words 3.18 [1.54] 4.01 [2.02] 0.83-0.83 [1.80] $0.012$
-Headings (H2–H4) per 1k words 0.92 [0.41] 0.84 [0.39] 0.08 [0.37] $0.075$
+Clean word count 14,200 [5,650] 9,360 [4,710] +4,840 $10^{-200}$
+Flesch–Kincaid grade 19.3 [6.6] 12.8 [9.9] +6.5 $10^{-100}$
+Lexical diversity (TTR) 0.244 [0.057] 0.248 [0.049] -0.004 0.02
+References per 1k words 21.4 [7.4] 90.2 [218] -68.8 $10^{-50}$
+Links per 1k words 16.8 [6.0] 428 [683] -411 $10^{-120}$
+Headings (H2–H4)/1k words 2.87 [1.20] 5.58 [8.32] -2.71 $10^{-40}$
 3.3 Cross-Platform Similarity
-We evaluated alignment between matched articles using lexical, semantic, structural, and stylistic metrics.
-Figure 3 presents the correlation matrix among these metrics.
-Lexical measures (TF–IDF, Jaccard, and $n$-gram overlap) tend to cluster with semantic similarity, while stylistic and structural measures form related but partially distinct dimensions.
-Table 2 reports the average values.
-Figure 3: Correlation among similarity metrics (Pearson’s $r$).
-Table 2: Average similarity metrics across matched pairs (mean [SD]).
+To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, structural, and stylistic similarity metrics for each article pair.
+Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while structural and stylistic dimensions vary more independently.
+Table 2 reports the average values across all pairs.
+Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,811 article pairs.
+Table 2: Average similarity metrics across matched pairs (N=1,811N{=}1{,}811).
 Metric Mean [SD]
-TF–IDF cosine 0.50 [0.22]
-Jaccard (unigram) 0.32 [0.16]
-Overlap (1-gram) 0.58 [0.23]
-Overlap (2-gram) 0.33 [0.34]
-Overlap (3-gram) 0.25 [0.37]
-Semantic cosine 0.73 [0.12]
-BERTScore (F1) 0.02 [0.24]
-Structural similarity 0.31 [0.26]
-Stylistic similarity 0.64 [0.08]
-3.4 Distributions of Similarity Scores
-To characterise the heterogeneity of alignment, Figure 4 shows the empirical distributions of all similarity scores across the matched pairs.
-Several metrics are unimodal and right-skewed (e.g., semantic cosine, BERTScore), suggesting generally high topical overlap.
-Others (e.g., higher-order $n$-gram overlap and structural similarity) exhibit broader or even multimodal shapes, consistent with clusters of near-verbatim matches versus paraphrastic or restructured articles.
-Figure 4: Distributions of article-to-article similarity scores across metrics.
-3.5 Does Length Drive Similarity?
-Finally, we test whether longer Wikipedia entries are more semantically aligned with GrokiPedia.
-Figure 5 plots semantic cosine against the (log) Wikipedia word count.
-The association is weak-to-moderate and not monotonic, indicating that content focus and selection likely matter more than sheer length for semantic convergence.
-Figure 5: Wikipedia length (log scale) vs. semantic similarity (cosine). A least-squares trend line is overlaid.
+TF–IDF cosine 0.505 [0.149]
+Jaccard (unigram) 0.362 [0.217]
+Overlap (1-gram) 0.597 [0.213]
+Overlap (2-gram) 0.329 [0.339]
+Overlap (3-gram) 0.250 [0.374]
+Semantic cosine 0.825 [0.106]
+BERTScore (F1) 0.193 [0.334]
+Structural similarity 0.216 [0.163]
+Stylistic similarity 0.835 [0.072]
+3.4 Distribution of Similarity Scores
+Figure 4 compiles the histograms for all nine similarity metrics.
+Most distributions are unimodal and right-skewed, especially lexical and semantic measures, indicating that the majority of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents.
+Structural similarity, by contrast, is more dispersed, reflecting Grokipedia’s simpler sectioning and reference patterns.
+Figure 4: Empirical distributions of similarity scores across metrics. Each panel corresponds to one of the nine metrics.
+Together, these findings reinforce that Grokipedia functions less as an independent encyclopedic corpus than as an AI-generated echo of Wikipedia, maintaining meaning and tone while attenuating the structural and evidentiary scaffolding that underpins human-curated knowledge.
 4 Discussion and Conclusion
-This study presents the first large-scale, systematic comparison between GrokiPedia and Wikipedia across 382 matched article pairs.
-Despite their distinct production paradigms—community-driven versus AI-generated—the two encyclopedias exhibit remarkable alignment at the level of meaning and style.
-Semantic similarity averages around 0.70, and stylistic similarity around 0.65, indicating that GrokiPedia reproduces much of Wikipedia’s linguistic and conceptual structure.
+This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 1,800 matched article pairs drawn from the 2,000 most-edited English Wikipedia titles.
+Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style.
+Semantic similarity averages around 0.82 and stylistic similarity around 0.83, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure.
 However, the platforms diverge substantially in form and informational scaffolding.
 Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels (greater syntactic complexity) but lower lexical diversity and reference density.
-This pattern suggests that GrokiPedia’s generation process elaborates on existing material—expanding text length and rhetorical flow—rather than producing substantively new or more rigorously sourced knowledge.
+This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
 The similarity distributions reveal further nuance.
-While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are bimodal, suggesting two distinct regimes of generation.
-In some cases, GrokiPedia closely mirrors Wikipedia’s organization and register; in others, it diverges sharply, producing verbose or restructured narratives.
+While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are more dispersed, suggesting heterogeneous modes of generation.
+Some Grokipedia pages closely mirror Wikipedia’s organization and tone; others diverge sharply, producing verbose or restructured narratives.
 This heterogeneity implies a hybrid authorship logic: part replication, part improvisation.
+These results also resonate with earlier linguistic analyses of Wikipedia itself.
+Yasseri, Kornai, and Kertész [2012a] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts.
+Our findings parallel this pattern in reverse, Grokipedia’s language inflates sentence length and readability grade without a proportional increase in lexical variety.
+In both cases, surface-level stylistic change (simplification or elaboration) alters perceived accessibility more than underlying vocabulary structure.
+Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
 Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [2, 7, 24].
-Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many GrokiPedia articles are derived from Wikipedia—often copied or paraphrased—while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21].
+Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21].
 Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density.
 In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution.
-Rather than “purging bias,” the system appears to re-encode it—subtly and opaquely—within model parameters and prompt conditioning.
+Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
 As argued by Yasseri [24], Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable.
 Grokipedia inverts this model.
-Its authorship is singular, automated, and invisible—bias becomes latent, hidden within model weights and unseen editorial heuristics.
+Its authorship is singular, automated, and invisible, bias becomes latent, hidden within model weights and unseen editorial heuristics.
 This creates what might be termed an epistemic opacity paradox: Grokipedia appears neutral because it lacks human editors, yet its underlying generative logic is uninspectable and unaccountable.
@@
 Several limitations qualify these findings.
-First, the dataset focuses on Wikipedia’s 416 most-edited pages, of which 382 have counterparts on GrokiPedia.
-This selection likely overrepresents high-profile, contentious topics—those most prone to editing wars or ideological scrutiny [25].
-Second, the similarity metrics employed—lexical, semantic, structural, and stylistic—quantify form and textual alignment but not factual accuracy or ideological framing.
+First, the dataset focuses on Wikipedia’s 2,000 most-edited pages, of which around 1,800 have counterparts on Grokipedia.
+This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [26].
+Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing.
 The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison.
@@
 It mirrors Wikipedia’s informational scope and linguistic tone but replaces community deliberation with algorithmic synthesis.
-The result is an encyclopedia that is fluent yet fragile—expansive in form but shallow in verifiability.
+The result is an encyclopedia that is fluent yet fragile, expansive in form but shallow in verifiability.
 In effect, Grokipedia trades collective accountability for computational authority.
@@
 As large language models increasingly mediate the production of knowledge, they risk conflating eloquence with accuracy and replacing transparency with automation.
-The challenge ahead is therefore not only technical—improving factuality or retrieval—but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge.
+The challenge ahead is therefore not only technical, improving factuality or retrieval, but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge.
 Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust.
-Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether GrokiPedia’s expansion of text corresponds to an expansion—or erosion—of understanding.
+Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether Grokipedia’s expansion of text corresponds to an expansion, or erosion, of understanding.
 5 Acknowledgement
-This publication has emanated from research supported in part by grants from Taighde Eireann – ´ Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.
+This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.

### v3

@@
 Abstract
-The launch of Grokipedia, an AI-generated encyclopedia developed by Elon Musk’s xAI, was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries via the large language model Grok.
+The launch of Grokipedia—an AI-generated encyclopedia developed by Elon Musk’s xAI—was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries using the Grok large language model.
 Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear.
-This study undertakes a large-scale computational comparison of 1,800 matched article pairs between Grokipedia and Wikipedia, drawn from the 2,000 most-edited Wikipedia pages.
-Using metrics across lexical richness, readability, structural organization, reference density, and semantic similarity, we assess how closely the two platforms align in form and substance.
-The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and greater structural variability.
-These findings suggest that AI-generated encyclopedic content currently mirrors Wikipedia’s informational scope but diverges in editorial norms, favoring narrative expansion over citation-based verification.
-The implications highlight new tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
+This study conducts a large-scale computational comparison of more than 17,000 matched article pairs from the 20,000 most-edited English Wikipedia pages.
+Using metrics spanning lexical richness, readability, reference density, structural features, and semantic similarity, we assess how closely the two platforms align in form and substance.
+We find that Grokipedia articles are substantially longer and contain significantly fewer references per word.
+Moreover, Grokipedia’s content divides into two distinct groups: one that remains semantically and stylistically aligned with Wikipedia, and another that diverges sharply.
+Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited news sources, concentrated primarily in entries related to politics, history, and religion.
+These findings suggest that AI-generated encyclopedic content diverges from established editorial norms—favouring narrative expansion over citation-based verification.
+The implications highlight emerging tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
 1 Introduction
@@
 Despite its success, Wikipedia has long faced questions about bias, reliability, and systemic underrepresentation.
-[^1] For a survey of evidence and debates on reliability and bias, see e.g., [4, 8, 17, 18].
+[^1] For a survey of evidence and debates on reliability and bias, see e.g., [GreensteinZhu2012AER, ReliabilityWikipedia2019, MIWikipediaBias2024, RightingWriters2025].
 In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia.
-[^2] See [20, 1, 19, 11, 13, 6, 10].
-According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[20, 11, 10] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800–900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[20, 19, 1] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[13, 6, 11, 10]
+[^2] See [GrokipediaWP, AP2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025].
+According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[GrokipediaWP, BI2025, AP2025] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[Verge2025, Wired2025, Guardian2025, Time2025]
 The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better?
-Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[4] topical and coverage imbalances,[8] and framing asymmetries across ideology and gender.[17, 18] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[26] Parallel work on large language models (LLMs) shows measurable political and cultural biases that vary across model families and prompts,[22, 15, 3, 9] raising distinct questions about the epistemic reliability of AI-generated text.
+Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[GreensteinZhu2012AER] topical and coverage imbalances,[ReliabilityWikipedia2019] and framing asymmetries across ideology and gender.[MIWikipediaBias2024, RightingWriters2025] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[yasseri2012dynamics]
+Parallel research on large language models (LLMs) demonstrates measurable political and cultural biases that vary across model architectures and prompts [YangMenczer2023, Rettenberger2024, Exler2025, Jenny2023].
+LLMs exhibit human-like biases based on input framing [lior2025wildframe] and show intrinsic political biases that affect text continuation and prediction [lin2024investigating], raising questions about the epistemic reliability of AI-generated content.
+Grokipedia, in particular, relies on the Grok LLM developed by xAI.
+In comparative studies on bibliographic reference retrieval, Grok stood out as one of the few models that did not fabricate references, although roughly 40% of references contained incomplete or inaccurate details, such as publication year, journal name, or author information [sidorenko2025assessing].
+Similarly, in another study, researchers evaluated four LLMs, including Grok 2, to examine how source framing influences model judgments.
+Their results indicate that Grok 2 demonstrates high consistency and minimal ideological bias, with only modest shifts in evaluation when the framing of the input changed [germani2025source].
+These results suggest that Grok may have the potential to generate reliable and neutral encyclopedic content, although intrinsic biases in LLMs could still persist, raising the question of whether Grok can produce a less biased encyclopedia than Wikipedia.
 Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice?
-We address this by pairing articles across platforms and comparing them along four analytic dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), (3) structural similarity (section and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, part-of-speech composition).
+We address this by pairing articles across platforms and comparing them along three similarity dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), and (3) stylistic similarity (readability, lexical diversity, part-of-speech composition).
 In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
-To construct a broad and representative corpus, we began with Wikipedia’s 2,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[26] Among these, approximately 1,800 titles had valid, nontrivial matches on Grokipedia.
-This enlarged sample enables more stable estimates of cross-platform differences than earlier pilot comparisons limited to a few hundred pairs.
+We also measure the political bias of each article by aggregating the bias scores of its cited references (see Methods for details).
+To contextualize these patterns, we classify all articles into topical domains and then analyse how both similarity and the shift in political bias vary across topics.
+To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[yasseri2012dynamics] Among these, approximately 17,700 titles had valid, nontrivial matches on Grokipedia.
 This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia.
@@
 By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI.
-[^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [11, 6, 10, 23].
+[^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, GrogConversation2025].
 2 Methods
 2.1 Sampling and Data Collection
-We analyzed the 2,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts.
-Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [26].
-This sampling strategy thus prioritizes articles that are both textually substantial and socially significant.
-For each title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title>
-and
+We analyzed the 20,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts.
+Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [yasseri2012dynamics].
+To ensure that our comparison focused on substantive articles, we excluded all list-style pages as well as titles that were date- or year-like rather than topical.
+Specifically, we removed (a) pages with the title ”List of Topic” (e.g., UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
+For each remaining title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title> and
 https://Grokipedia.com/page/<Title>.
-HTML pages were downloaded via the requests library (Python 3.11) with polite delays and a standard user-agent header, then parsed using BeautifulSoup4.
-The extraction and analysis pipeline continuously wrote progress files (wiki_Grokipedia_similarity_live.csv) to enable incremental recovery and monitoring.
-After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained.
-Of the 2,000 target titles, 1,811 matched article pairs satisfied these criteria and formed the final analytical sample.
+HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11) with polite delays and a standard user-agent header, and then parsed using BeautifulSoup4.
+After host-aware text extraction (described below), we retained only article pairs in which both platforms produced at least 500 words of clean prose.
+Of the original 20,000 target titles, 17,790 matched pairs met these criteria and formed the final analytical sample.
 2.2 Text Cleaning and Feature Extraction
-We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
+We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximise content fidelity.
 For Wikipedia, extraction was restricted to the #mw-content-text container and limited strictly to <p> and <li> elements.
@@
 Scripts, styles, and sidebars were removed in both cases.
-Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, structural, and stylistic metrics was computed:
-  • Structural features: counts of paragraphs, headings (h1–h4), hyperlinks, images, and references; visible character counts; and derived ratios such as references, links, and headings per 1,000 words.
+Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, and stylistic metrics was computed.
   • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
@@
 2.3 Similarity Measures Between Platforms
-To quantify cross-platform alignment, we computed a suite of nine similarity measures grouped into four conceptual domains:
+To quantify cross-platform alignment, we computed a suite of eight similarity measures grouped into four conceptual domains:
   1. 1.
@@
   3. 3.
-    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14], and contextual similarity via BERTScore F1 for the first 50 sentences per article [27].
+    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[reimers2019sentence], and contextual similarity via BERTScore F1 for the first 50 sentences per article [zhang2019bertscore].
   4. 4.
-    Structural and stylistic similarity: a composite of (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths (weighted 0.6 and 0.4, respectively), together with normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
-A composite similarity index was calculated as the mean of lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
+    stylistic similarity: The Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
+To verify the extent to which these metrics reflected a single latent construct, we conducted a principal components analysis (PCA) on the full set of similarity measures.
+The first principal component accounted for the vast majority of total variance, with all measures loading in the expected (positive) direction.
+This indicated that the metrics were largely co-linear and captured a common underlying dimension of article similarity.
+Based on this result, we constructed a unified combined similarity score, defined as the standardized value of the first principal component, which we use in all subsequent analyses.
+2.4 Political Bias ratings
+Outlet-level political-leaning scores were taken from the News Media Bias and Factuality dataset [sanchez2024mapping], mapped to a symmetric numeric scale, and used only when a numeric bias was available for a domain or its brand variant (e.g., bbc.com/bbc.co.uk).
+Non-news and reference/utility hosts (e.g., doi.org, archive.org, worldcat.org, major platforms, and *.gov/*.edu) were excluded.
+For brand merging we used eTLD+1 normalization plus a brand key (leftmost label).
+For each article, bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift).
+Centers of mass in Figure 6 were derived using Gaussian kernel density estimation (KDE) applied to the raw article-level points, computed separately for the low- and high-similarity partitions.
+For the pie charts in Figure 5, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 40% cumulative share of referenced news outlets.
+Slice colors encode outlet bias, and panel titles report the mention-weighted average platform bias normalized to ±100%.
+2.5 Topics
+To categorize the topics of the articles, we developed a multi-stage large-language-model (LLM) classification pipeline.
+We first constructed a fixed taxonomy of fourteen semantically distinct categories (e.g., History, Politics and Political Figures, Science and Technology, Literature and Art).
+Article titles were processed in batches of up to 300 and passed to the GPT-5-nano-2025-08-07 model with strict instructions to assign each title to exactly one category and to avoid the fallback use of “Other” whenever a reasonable category was available.
+The model returned a JSON-formatted list of titles grouped by category, and the script automatically retried up to five times if the output contained formatting errors.
+All valid batch outputs were then combined into an initial classification table mapping each title to its assigned category.
+Because LLM-generated labels can occasionally be inconsistent across runs, we used a multi-round refinement procedure to improve reliability.
+The entire classification process was repeated across several independent rounds using the same model and prompt.
+After each round, titles labeled “Other” were separated and given to the model for reclassification.
+We then merged the rounds.
+Any article still lacking a category after all rounds was assigned “Other”.
+This multi-round integration approach increased coverage, reduced noise, and yielded a more stable and robust set of topic labels for all article pairs.
+Articles were assigned to one of the following categories:
+  • History
+  • Politics and Political Figures
+  • Geography, Nations, and Cities
+  • Music, Movies, and Entertainment Celebrities
+  • Sports, Athletes, and Teams
+  • Science, Technology, and Knowledge
+  • Religion and Ideologies
+  • Brands and Products
+  • Languages
+  • Literature and Art
+  • Health and Environment
+  • Business and Infrastructure
+  • Animals and Nature
+  • Other
 2.6 Statistical and Comparative Analysis
@@
 Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering.
-Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib and exported directly to PDF for inclusion in the manuscript.
+Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib.
 2.7 Limitations
 The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia.
-Because edit frequency correlates with controversy and visibility [26], the sample likely overrepresents politically or culturally salient topics.
+Because edit frequency correlates with controversy and visibility [yasseri2012dynamics], the sample likely overrepresents politically or culturally salient topics.
 In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose.
@@
 3.1 Sample and Coverage
-From the expanded list of the 2,000 most-edited English Wikipedia titles, we successfully retrieved and matched 1,811 Grokipedia–Wikipedia article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each).
-[^4] Clean word counts were computed from visible prose only, excluding infoboxes, scripts, and navigation templates.
-This larger corpus substantially extends the coverage of the original 416-title sample while preserving topical breadth and diversity.
+From the list of the 20,000 most-edited English Wikipedia articles, we successfully retrieved and matched 17,790 Grokipedia–Wikipedia article pairs that passed all quality checks (see more details in Methods).
+This corpus provides substantial coverage across a wide and diverse range of topics.
 3.2 Descriptive Differences Between Platforms
-Figure 1 compares article-length distributions on a logarithmic scale.
-Grokipedia entries remain systematically longer than their Wikipedia counterparts, though the gap narrows compared with the smaller sample.
-Readability, measured by the Flesch–Kincaid grade level (panel B of Figure 2), is again higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
+Figure 1 compares the distribution of article lengths across platforms on a logarithmic scale.
+Overall, Grokipedia entries are systematically longer than their Wikipedia counterparts.
+While Wikipedia contains a larger number of very short articles, Grokipedia articles exhibit a pronounced peak around 7,000 words, indicating that most Grokipedia entries are substantially more verbose.
+Readability, measured by the Flesch–Kincaid grade level (Figure 2 (b)) and lexical diversity (Figure 2 (c), is higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
 Table 1 summarises key descriptive statistics.
-Across the enlarged dataset, Grokipedia articles average about 14,200 words versus 9,400 on Wikipedia, and they exhibit lower lexical diversity and far fewer explicit references, links, and headings per thousand words.
+Across the dataset, Grokipedia articles average 7,662 words versus 6,280 on Wikipedia, and they exhibit far fewer explicit references, links, and headings per thousand words.
 These contrasts reinforce that Grokipedia expands text length through elaboration rather than citation density or structural detail.
@@
 Panels show: (a) clean word count (log10), (b) Flesch–Kincaid grade, (c) lexical diversity (TTR), (d) references per 1k words, (e) links per 1k words, and (f) headings (H2–H4) per 1k words.
-Table 1: Descriptive statistics across matched article pairs (N=1,811N{=}1{,}811). Values are means [SD]; $p$ from paired $t$-tests.
+Table 1: Descriptive differences between Grokipedia and Wikipedia articles.
+Means and standard deviations (in brackets) are reported for each measure.
+$\Delta$ denotes the mean difference (Grokipedia - Wikipedia), and $p$ reports two-sided paired comparison $p$-values
 Measure Grokipedia Wikipedia $\Delta$ $p$
-Clean word count 14,200 [5,650] 9,360 [4,710] +4,840 $10^{-200}$
-Flesch–Kincaid grade 19.3 [6.6] 12.8 [9.9] +6.5 $10^{-100}$
-Lexical diversity (TTR) 0.244 [0.057] 0.248 [0.049] -0.004 0.02
-References per 1k words 21.4 [7.4] 90.2 [218] -68.8 $10^{-50}$
-Links per 1k words 16.8 [6.0] 428 [683] -411 $10^{-120}$
-Headings (H2–H4)/1k words 2.87 [1.20] 5.58 [8.32] -2.71 $10^{-40}$
+Clean word count 7662.36 [3662.22] 6280.51 [3828.40] 1381.846 $10^{-10}$
+Flesch–Kincaid grade 14.53 [3.93] 10.67 [1.69] 3.859 $10^{-10}$
+Lexical diversity (TTR) 0.29 [0.05] 0.28 [0.06] 0.006 $10^{-10}$
+References per 1k words 20.78 [9.86] 35.03 [37.06] -14.257 $10^{-10}$
+Links per 1k words 24.18 [16.14] 356.58 [323.37] -332.405 $10^{-10}$
+Headings (H2–H4)/1k words 4.52 [1.97] 6.17 [5.19] -1.653 $10^{-10}$
 3.3 Cross-Platform Similarity
-To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, structural, and stylistic similarity metrics for each article pair.
-Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while structural and stylistic dimensions vary more independently.
+To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, and stylistic similarity metrics for each article pair.
+Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while stylistic dimensions vary more independently.
 Table 2 reports the average values across all pairs.
-Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,811 article pairs.
-Table 2: Average similarity metrics across matched pairs (N=1,811N{=}1{,}811).
-Metric Mean [SD]
-TF–IDF cosine 0.505 [0.149]
-Jaccard (unigram) 0.362 [0.217]
-Overlap (1-gram) 0.597 [0.213]
-Overlap (2-gram) 0.329 [0.339]
-Overlap (3-gram) 0.250 [0.374]
-Semantic cosine 0.825 [0.106]
-BERTScore (F1) 0.193 [0.334]
-Structural similarity 0.216 [0.163]
-Stylistic similarity 0.835 [0.072]
+Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,8287 article pairs.
+Table 2: Means and standard deviations (in brackets) are listed for each similarity metric, along with observed minimum and maximum values.
+Higher scores indicate greater similarity between Grokipedia and Wikipedia articles.
+(N=17,790N{=}17{,}790)
+Metric Mean [SD] Min Max
+TF–IDF cosine 0.684 [0.186] 0.024 1.000
+Jaccard (unigram) 0.434 [0.299] 0.046 1.000
+Overlap (1-gram) 0.640 [0.235] 0.190 1.000
+Overlap (2-gram) 0.411 [0.373] 0.048 1.000
+Overlap (3-gram) 0.373 [0.412] 0.005 1.000
+Semantic cosine 0.862 [0.102] 0.160 1.000
+BERTScore (F1) 0.306 [0.369] -0.606 0.983
+Stylistic similarity 0.882 [0.076] 0.429 0.999
 3.4 Distribution of Similarity Scores
-Figure 4 compiles the histograms for all nine similarity metrics.
-Most distributions are unimodal and right-skewed, especially lexical and semantic measures, indicating that the majority of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents.
-Structural similarity, by contrast, is more dispersed, reflecting Grokipedia’s simpler sectioning and reference patterns.
-Figure 4: Empirical distributions of similarity scores across metrics. Each panel corresponds to one of the nine metrics.
-Together, these findings reinforce that Grokipedia functions less as an independent encyclopedic corpus than as an AI-generated echo of Wikipedia, maintaining meaning and tone while attenuating the structural and evidentiary scaffolding that underpins human-curated knowledge.
+Figure 4 compiles the histograms for all eight similarity metrics.
+Most distributions are binomial, indicating that a portion of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents while the others are significantly distinct.
+Figure 4: Empirical distributions of similarity scores across metrics.
+Each panel corresponds to one of the eight metrics.
+(a) TF–IDF cosine, (b) Jaccard (unigram) (c) Overlap (1-gram) (d) Overlap (2-gram) (e) Semantic cosine (g) BERTScore (F1) (f) Stylistic similarity
+To reduce this multidimensional space, we conducted a principal components analysis (PCA).
+The first component explained the vast majority of variance, with all metrics loading in the expected direction, indicating that the measures capture a common underlying dimension of article similarity.
+We therefore use this component as a combined similarity score in subsequent analyses.
+Figure 5 shows the distribution of the combined similarity score.
+Figure 5: The distribution of the combined similarity scores.
+The distribution in Figure 5 is distinctly bimodal, suggesting the presence of two substantive groups of article pairs: one in which Grokipedia and Wikipedia differ substantially, and another in which the two versions are highly similar.
+3.5 Political Bias
+Another dimension of article change concerns shifts in political orientation.
+To quantify this, we examined the references cited in each article and assigned each source a bias score using a publicly available dataset (see Methods for details).
+For each article pair, we computed the bias shift as the difference between the average bias score of the Grokipedia article and that of its Wikipedia counterpart.
+Positive values indicate that the Grokipedia version relies more heavily on right-leaning sources, whereas negative values reflect a shift toward left-leaning sources.
+Figure 6 plots the bias shift for each article pair against its combined similarity score.
+As shown in the figure, articles that differ more substantially across platforms are markedly more likely to exhibit a right-leaning shift, whereas highly similar articles tend to cluster near zero, indicating minimal ideological movement.
+Figure 6: Bias shift vs.
+article similarity.
+Each point represents an article pair.
+The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions (higher values indicate greater similarity), and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia).
+Positive values indicate a shift toward right-leaning sources, while negative values indicate a shift toward left-leaning sources.
+Colored markers denote the centres of mass for the low-similarity and high-similarity groups.
+As similarity increases, bias shifts converge toward zero; when similarity is low, the magnitude and dispersion of bias shifts increase.
+To further examine differences in the sources cited within the least similar article pairs, we focus on the subset of articles with a negative combined similarity score.
+Figure 5 displays the top 40% most frequently cited news sources within this low-similarity group for each platform.
+Each wedge is colour-coded according to the ideological bias of the outlet.
+As shown in the figure, the weighted average bias score of Grokipedia articles in this subset is approximately 25% left-leaning, compared with 31% left-leaning for Wikipedia.
+This difference indicates a relative rightward shift in the sourcing of the least similar Grokipedia articles, consistent with the pattern observed in the bias–similarity relationship above.
+Figure 7: Top news sources among low-similarity articles (negative combined similarity score).
+Wedges show each outlet’s share of all mentions within this subset; colors reflect the outlet’s political-leaning score from the public News Media Bias and Factuality dataset.
+[^5] https://huggingface.co/datasets/sergioburdisso/news_media_bias_and_factuality Titles report the overall (news-only) mention-weighted bias for each platform (normalized to $\pm 100\%$).
+3.6 Comparison within Topics
+Because these articles cover a diverse set of subject areas, we additionally classified each article using GPT-5 (see Methods for details).
+This allows us to compare similarity and bias shift not only across individual articles but also across broader thematic domains.
+Figure 8 summarizes these findings by plotting, for each topical category, the average similarity between Wikipedia and Grokipedia articles together with the corresponding average shift in political bias.
+As shown in Figure 8(a), the largest cross-platform differences in similarity appear in articles related to politics, geography, history, business, and religion.
+In parallel, as shown in Figure 8(b), the strongest rightward shifts in source bias occur in articles on religion, history, languages, and business, indicating that ideological divergence is especially pronounced in these domains.
+Figure 8: Topic-level comparison of similarity and political bias shift between Wikipedia and Grokipedia articles.
+(a) Mean combined similarity scores between Wikipedia and Grokipedia articles within each topic, with standard-error bars.
+(b) Mean political bias shift (Grokipedia – Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia.
+Error bars show standard errors within each topic.
 4 Discussion and Conclusion
-This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 1,800 matched article pairs drawn from the 2,000 most-edited English Wikipedia titles.
+This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 17,700 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles.
 Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style.
-Semantic similarity averages around 0.82 and stylistic similarity around 0.83, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure.
+Semantic similarity averages around 0.86 and stylistic similarity around 0.88, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure.
 However, the platforms diverge substantially in form and informational scaffolding.
-Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels (greater syntactic complexity) but lower lexical diversity and reference density.
+Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels and lexical diversity (greater syntactic complexity), but reference density.
 This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
-The similarity distributions reveal further nuance.
-While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are more dispersed, suggesting heterogeneous modes of generation.
-Some Grokipedia pages closely mirror Wikipedia’s organization and tone; others diverge sharply, producing verbose or restructured narratives.
-This heterogeneity implies a hybrid authorship logic: part replication, part improvisation.
+An important distinction between the two platforms lies in their use of supporting media and internal link structure.
+Unlike Wikipedia, the current version of Grokipedia includes no images and does not offer linking to related pages within the platform.
+These omissions substantially reduce navigational affordances and make Grokipedia articles harder to scan, explore, and contextualize.
+The similarity distributions reveal additional nuance.
+The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply.
+This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation.
+In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it departs substantially from the source material, generating content that is stylistically and semantically distinct.
 These results also resonate with earlier linguistic analyses of Wikipedia itself.
-Yasseri, Kornai, and Kertész [2012a] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts.
-Our findings parallel this pattern in reverse, Grokipedia’s language inflates sentence length and readability grade without a proportional increase in lexical variety.
+Yasseri, Kornai, and Kertész [yasseri2012complexity] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts.
+Our findings parallel this pattern in reverse: Grokipedia’s language inflates sentence length and readability grade.
 In both cases, surface-level stylistic change (simplification or elaboration) alters perceived accessibility more than underlying vocabulary structure.
 Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
-Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [2, 7, 24].
-Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21].
+Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [apnews2025, businessinsider2025, yasseri2025conversation].
+Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025].
 Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density.
@@
 Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
-As argued by Yasseri [24], Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable.
+As argued by yasseri2025conversation, Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable.
 Grokipedia inverts this model.
@@
 The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
+Complementing these textual and stylistic findings, our reference-level analysis (Figures 65) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity.
+When Grokipedia reproduces Wikipedia closely (high similarity), the resulting entries show little or no bias shift.
+However, low-similarity articles—those most substantially rewritten—display both greater variance and a systematic rightward tendency in bias shift.
+This pattern suggests that editorial divergence in Grokipedia is concentrated precisely where generative rewriting occurs.
+The corresponding breakdown of reference sources further supports this interpretation.
+In the low-similarity subset, citation networks are dominated by a narrower set of news outlets with identifiable political leanings, as measured by the News Media Bias and Factuality dataset [sanchez2024mapping].
+Whereas Wikipedia’s reference base remains diverse and balanced, Grokipedia’s alternative sourcing emphasizes ideologically marked or partisan media domains.
+Together, these results indicate that Grokipedia’s generative process may amplify latent source bias when departing from human-authored baselines, substituting breadth and attribution with ideologically skewed curation.
+The topic-level analysis reveals that cross-platform divergence is not evenly distributed but concentrated in specific domains.
+Articles on politics, geography, history, business, and religion show the lowest similarity between Grokipedia and Wikipedia, suggesting that the model’s generative behavior departs most from human-edited content precisely in areas where public narratives are contested or where factual structures are complex.
+Notably, some of these domains exhibit the strongest rightward shift in cited sources, too.
+This alignment between low similarity and increased ideological skew suggests that Grokipedia’s deviations are not merely stylistic but often substantively directional.
 Several limitations qualify these findings.
-First, the dataset focuses on Wikipedia’s 2,000 most-edited pages, of which around 1,800 have counterparts on Grokipedia.
-This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [26].
+First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which around 17,700 have counterparts on Grokipedia.
+This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [yasseri2012dynamics].
 Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing.

### v4

@@
 Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear.
-This study conducts a large-scale computational comparison of more than 17,000 matched article pairs from the 20,000 most-edited English Wikipedia pages.
+This study conducts a large-scale computational comparison of 17,790 matched article pairs from the 20,000 most-edited English Wikipedia pages.
 Using metrics spanning lexical richness, readability, reference density, structural features, and semantic similarity, we assess how closely the two platforms align in form and substance.
@@
 Moreover, Grokipedia’s content divides into two distinct groups: one that remains semantically and stylistically aligned with Wikipedia, and another that diverges sharply.
-Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited news sources, concentrated primarily in entries related to politics, history, and religion.
-These findings suggest that AI-generated encyclopedic content diverges from established editorial norms—favouring narrative expansion over citation-based verification.
-The implications highlight emerging tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
+Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited sources, concentrated primarily in entries related to politics, history, and religion.
+More broadly, the findings indicate that AI-generated encyclopedic content departs from established editorial norms, favoring narrative expansion over citation-based verification, raising questions about transparency, provenance, and the governance of knowledge in automated information systems.
 1 Introduction
@@
 In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia.
-[^2] See [GrokipediaWP, AP2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025].
-According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[GrokipediaWP, BI2025, AP2025] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[Verge2025, Wired2025, Guardian2025, Time2025]
+[^2] See [GrokipediaWP, apnews2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025].
+According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance [GrokipediaWP, BI2025, apnews2025].
+Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models [Verge2025, Wired2025, Guardian2025, Time2025].
 The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better?
@@
 These results suggest that Grok may have the potential to generate reliable and neutral encyclopedic content, although intrinsic biases in LLMs could still persist, raising the question of whether Grok can produce a less biased encyclopedia than Wikipedia.
-Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice?
+Against this backdrop, our study asks a concrete, testable question: for titles that exist on both platforms, how similar are Grokipedia and Wikipedia in practice?
 We address this by pairing articles across platforms and comparing them along three similarity dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), and (3) stylistic similarity (readability, lexical diversity, part-of-speech composition).
 In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
-We also measure the political bias of each article by aggregating the bias scores of its cited references (see Methods for details).
+We also measure the political bias of each article by aggregating the bias scores of its cited sources (see Methods for details).
 To contextualize these patterns, we classify all articles into topical domains and then analyse how both similarity and the shift in political bias vary across topics.
-To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[yasseri2012dynamics] Among these, approximately 17,700 titles had valid, nontrivial matches on Grokipedia.
+To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience [yasseri2012dynamics].
+Among these, 17,790 titles had valid, nontrivial matches on Grokipedia.
 This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia.
@@
 Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability.
-By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI.
-[^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, GrogConversation2025].
+By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI  [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, yasseri2025conversation]..
 2 Methods
@@
 We analyzed the 20,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts.
-Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [yasseri2012dynamics].
 To ensure that our comparison focused on substantive articles, we excluded all list-style pages as well as titles that were date- or year-like rather than topical.
-Specifically, we removed (a) pages with the title ”List of Topic” (e.g., UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
+Specifically, we removed (a) pages with the title ”List of Topic” (e.g., List of UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
 For each remaining title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title> and
 https://Grokipedia.com/page/<Title>.
-HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11) with polite delays and a standard user-agent header, and then parsed using BeautifulSoup4.
+HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11), and then parsed using BeautifulSoup4.
 After host-aware text extraction (described below), we retained only article pairs in which both platforms produced at least 500 words of clean prose.
 Of the original 20,000 target titles, 17,790 matched pairs met these criteria and formed the final analytical sample.
-2.2 Text Cleaning and Feature Extraction
-We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximise content fidelity.
-For Wikipedia, extraction was restricted to the #mw-content-text container and limited strictly to <p> and <li> elements.
-Elements such as infoboxes, hatnotes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
-For Grokipedia, a more adaptive extractor identified the main article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high link-density regions.
-Scripts, styles, and sidebars were removed in both cases.
-Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, and stylistic metrics was computed.
-  • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
-  • Lexical density and reading time: ratio of unique to total words and estimated reading time assuming 200 words per minute.
+2.1.1 Host-aware text and feature extraction
+We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
+For Wikipedia, extraction was restricted to the #mw-content-text container and limited to <p> and <li> elements.
+Infoboxes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
+For Grokipedia, we used a more adaptive extractor to identify the primary article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high–link-density regions.
+Scripts, style elements, and sidebars were removed for both platforms.
+We tokenized each cleaned article into sentences and words using nltk’s Punkt tokenizer.
+From the extracted text of each article, we computed a set of stylistic and readability metrics, in addition to lexical density and estimated reading time.
+All features were calculated separately for each platform.
+  • Average sentence length: Computed as the mean number of tokens per sentence in the extracted text.
+This measure captures syntactic complexity, with longer sentences generally reflecting more complex or formally structured writing.
+  • Lexical diversity (type–token ratio): Defined as the number of unique word types divided by the total number of word tokens.
+Higher values indicate richer vocabulary usage and lower repetition within the text.
+  • Flesch–Kincaid grade level: A widely used readability formula that estimates the U.S.
+school grade level required to understand the text.
+It combines average sentence length and average syllables per word, with higher scores indicating more difficult material.
+  • Gunning–Fog index: A readability metric evaluating the number of complex words (three or more syllables) and sentence length.
+It approximates the years of formal education needed to comprehend a text on first reading and is commonly used to quantify formal, technical, or academic writing.
+  • Part-of-speech (POS) composition: Using spaCy’s en_core_web_sm tagger, we computed the proportion of nouns, verbs, adjectives, and adverbs in each article.
+POS distribution provides a coarse-grained linguistic profile of stylistic variation, such as noun-dense factual exposition versus verb-heavy narrative text.
+  • Lexical density: Calculated as the ratio of unique words (types) to total words (tokens).
+Lexical density reflects how information-dense a text is; higher values indicate a more compact and content-rich lexical profile.
+  • Estimated reading time: Derived by dividing the total word count by a constant reading rate of 200 words per minute.
+This measure offers a standardized estimate of the amount of content a typical reader must process.
 All features were computed independently for both platforms and labeled with prefixes a_ (Grokipedia) and b_ (Wikipedia).
 2.2 Similarity Measures Between Platforms
-To quantify cross-platform alignment, we computed a suite of eight similarity measures grouped into four conceptual domains:
+To quantify alignment between articles on the two platforms, we computed four classes of similarity measures, each capturing a distinct dimension of textual resemblance:
   1. 1.
-    Lexical similarity: cosine similarity of TF–IDF vectors (1–2 grams, English stop-words removed) and unigram Jaccard index.
+    Lexical similarity: Lexical similarity was assessed using two complementary approaches.
+First, we computed cosine similarity between TF–IDF vectors constructed from 1–2 gram features with standard English stop-words removed.
+This measure captures similarity in word- and phrase-level vocabulary usage weighted by informativeness.
+Second, we calculated the unigram Jaccard index, defined as the ratio of the intersection to the union of unique word sets across platforms.
+This index evaluates the extent to which the two texts draw from the same vocabulary.
   2. 2.
-    N-gram overlap: overlap coefficients for 1-, 2-, and 3-gram sequences, capturing local phrase reuse.
+    N-gram overlap: To capture local phrase reuse and short-range structural similarity, we computed overlap coefficients for 1-, 2-, and 3-gram sequences.
+Each coefficient is defined as the size of the intersection of n-grams divided by the smaller of the two n-gram sets, providing a symmetric measure of shared phrase content independent of text length.
   3. 3.
-    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[reimers2019sentence], and contextual similarity via BERTScore F1 for the first 50 sentences per article [zhang2019bertscore].
+    Semantic similarity: We quantified semantic alignment using two embedding-based approaches. Cosine similarity between SentenceTransformer embeddings
+    (all-MiniLM-L6-v2) measured the overall semantic closeness of the full articles in a dense representation space.
+In addition, we computed contextual similarity using BERTScore F1 on the first 50 sentences of each article, which evaluates token-level semantic correspondence using contextual embeddings [reimers2019sentencebert].
+Together, these metrics assess similarity in meaning beyond surface-level wording.
   4. 4.
-    stylistic similarity: The Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
+    Stylistic similarity: Stylistic similarity was evaluated by computing the Manhattan distance between the two articles’ stylistic feature vectors, which included sentence length, lexical diversity, readability metrics, and part-of-speech composition.
+The resulting distances were transformed linearly to a 0–1 scale, where higher values indicate stronger similarity.
+This measure captures convergence in writing style independent of content.
 To verify the extent to which these metrics reflected a single latent construct, we conducted a principal components analysis (PCA) on the full set of similarity measures.
@@
 2.3 Political Bias ratings
-Outlet-level political-leaning scores were taken from the News Media Bias and Factuality dataset [sanchez2024mapping], mapped to a symmetric numeric scale, and used only when a numeric bias was available for a domain or its brand variant (e.g., bbc.com/bbc.co.uk).
-Non-news and reference/utility hosts (e.g., doi.org, archive.org, worldcat.org, major platforms, and *.gov/*.edu) were excluded.
-For brand merging we used eTLD+1 normalization plus a brand key (leftmost label).
-For each article, bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift).
-Centers of mass in Figure 6 were derived using Gaussian kernel density estimation (KDE) applied to the raw article-level points, computed separately for the low- and high-similarity partitions.
-For the pie charts in Figure 5, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 40% cumulative share of referenced news outlets.
-Slice colors encode outlet bias, and panel titles report the mention-weighted average platform bias normalized to ±100%.
-2.5 Topics
-To categorize the topics of the articles, we developed a multi-stage large-language-model (LLM) classification pipeline.
-We first constructed a fixed taxonomy of fourteen semantically distinct categories (e.g., History, Politics and Political Figures, Science and Technology, Literature and Art).
-Article titles were processed in batches of up to 300 and passed to the GPT-5-nano-2025-08-07 model with strict instructions to assign each title to exactly one category and to avoid the fallback use of “Other” whenever a reasonable category was available.
-The model returned a JSON-formatted list of titles grouped by category, and the script automatically retried up to five times if the output contained formatting errors.
-All valid batch outputs were then combined into an initial classification table mapping each title to its assigned category.
-Because LLM-generated labels can occasionally be inconsistent across runs, we used a multi-round refinement procedure to improve reliability.
-The entire classification process was repeated across several independent rounds using the same model and prompt.
-After each round, titles labeled “Other” were separated and given to the model for reclassification.
-We then merged the rounds.
-Any article still lacking a category after all rounds was assigned “Other”.
-This multi-round integration approach increased coverage, reduced noise, and yielded a more stable and robust set of topic labels for all article pairs.
-Articles were assigned to one of the following categories:
-  • History
-  • Politics and Political Figures
-  • Geography, Nations, and Cities
-  • Music, Movies, and Entertainment Celebrities
-  • Sports, Athletes, and Teams
-  • Science, Technology, and Knowledge
-  • Religion and Ideologies
-  • Brands and Products
-  • Languages
-  • Literature and Art
-  • Health and Environment
-  • Business and Infrastructure
-  • Animals and Nature
-  • Other
+We extracted all outbound hyperlinks contained in each article and mapped their host domains to political-leaning scores using the News Media Bias and Factuality dataset [sanchez2024mapping].
+Scores from this dataset were converted to a symmetric numeric scale (-1:left to 1:right) and were included only when a quantitative bias value was available for the domain or any of its recognized brand variants (e.g., bbc.com/bbc.co.uk).
+To consolidate hosts referring to the same outlet, we applied eTLD+1 normalization supplemented by a brand key derived from the leftmost label.
+This procedure ensured consistent attribution across domain variants and subdomain structures.
+The bias score of each article is the weighted average of the bias score of each domain based on the number of times the domain is mentioned in the page.
+For each article pair, the bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift).
+The centres of mass shown in Figure 1 were obtained via Gaussian kernel density estimation (KDE) applied to the article-level scatter points, computed separately for the low- and high-similarity partitions.
+For the pie charts in Figure 1, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 5 of the cumulative share of referenced news outlets.
+In this analysis, we focused exclusively on news sources, removing reference or utility hosts (e.g., doi.org, archive.org, worldcat.org), major platforms, and government or educational domains (.gov, .edu).
+2.4 Categorisation
+To assign each article to a topical domain, we implemented a multi-stage large-language-model (LLM) classification pipeline.
+Article titles were processed in batches of up to 300 and submitted to the gpt-5-nano-2025-08-07 model with explicit instructions to assign each title to exactly one topic and to avoid the use of “Other” unless no reasonable category applied.
+The model returned a JSON-formatted list of titles grouped by category.
+Outputs with formatting errors were automatically resubmitted, with up to five retries per batch.
+All valid outputs were merged into an initial table mapping each title to a category.
+Because LLM-generated labels can vary across runs, we conducted several independent classification rounds using the same model and prompt.
+After each round, titles labeled as “Other” were isolated and resubmitted for reclassification.
+The results were then merged across rounds.
+Titles that remained unclassified after all iterations were ultimately assigned to “Other”.
+This multi-round integration procedure increased coverage, reduced inconsistency, and yielded a more stable set of topic labels for all article pairs.
+As a final validation step, we submitted each provisional category (batched in sets of 200 entries) to the gpt-4.1-mini model and instructed it to flag and propose revised labels for any titles that appeared misclassified.
+All suggested corrections were manually reviewed before inclusion in the final categorisation.
+Articles were assigned to one of the following topical domains: history, politics, geography, entertainment, sports, science and technology, religion and ideologies, brands and products, languages, literature and art, health and environment, business and infrastructure, animals and nature, or other.
 2.5 Statistical and Comparative Analysis
 All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn.
-Descriptive statistics (mean $\pm$ SD) were calculated for all features.
-Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise.
+We calculated the descriptive statistics (mean $\pm$ SD) for all features.
+We evaluated Differences between Grokipedia and Wikipedia using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise.
 Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering.
@@
 2.6 Limitations
-The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia.
-Because edit frequency correlates with controversy and visibility [yasseri2012dynamics], the sample likely overrepresents politically or culturally salient topics.
-In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose.
-Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
+Our dataset is affected by several limitations.
+First, the dataset is drawn from Wikipedia’s 20,000 most-edited English-language pages, of which 17,790 have counterparts on Grokipedia.
+This selection likely overrepresents high-profile and contentious topics—those most prone to editing disputes and ideological scrutiny [yasseri2012dynamics].
+Second, the similarity metrics employed—lexical, semantic, structural, and stylistic—capture textual form and alignment but do not assess factual accuracy or ideological framing.
+As a result, hallucinated claims, selective omissions, or more subtle rhetorical biases remain beyond the scope of automated comparison.
+Third, the bias measure is based on the ideological orientation of source domains rather than individual articles, relies on a limited dataset, and is grounded in the U.S.
+political context, which may not generalize to other media systems.
+Fourth, both platforms are dynamic: Grokipedia and Wikipedia continue to evolve, and our analysis reflects only a snapshot of their content at the time of data collection.
+Finally, Grokipedia’s underlying training data, retrieval mechanisms, and editorial interventions remain opaque, precluding full provenance auditing and limiting causal inference about the sources of observed bias.
 3 Results
@@
 Each point represents an article pair.
-The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions (higher values indicate greater similarity), and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia).
+The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions, and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia).
 Positive values indicate a shift toward right-leaning sources, while negative values indicate a shift toward left-leaning sources.
-Colored markers denote the centres of mass for the low-similarity and high-similarity groups.
-As similarity increases, bias shifts converge toward zero; when similarity is low, the magnitude and dispersion of bias shifts increase.
-To further examine differences in the sources cited within the least similar article pairs, we focus on the subset of articles with a negative combined similarity score.
-Figure 5 displays the top 40% most frequently cited news sources within this low-similarity group for each platform.
-Each wedge is colour-coded according to the ideological bias of the outlet.
-As shown in the figure, the weighted average bias score of Grokipedia articles in this subset is approximately 25% left-leaning, compared with 31% left-leaning for Wikipedia.
-This difference indicates a relative rightward shift in the sourcing of the least similar Grokipedia articles, consistent with the pattern observed in the bias–similarity relationship above.
-Figure 7: Top news sources among low-similarity articles (negative combined similarity score).
-Wedges show each outlet’s share of all mentions within this subset; colors reflect the outlet’s political-leaning score from the public News Media Bias and Factuality dataset.
-[^5] https://huggingface.co/datasets/sergioburdisso/news_media_bias_and_factuality Titles report the overall (news-only) mention-weighted bias for each platform (normalized to $\pm 100\%$).
+Colored markers denote the centres of mass for the low- and high-similarity groups.
+Inset: Pie charts show the top five news domains cited in Grokipedia (left) and Wikipedia (right) among articles with negative combined similarity scores, weighted by citation frequency.
+Slice areas represent relative contribution, and colors encode political bias.
+The inset of Figure 6 displays the five most frequently cited news sources within this low-similarity group for each platform, with wedges color-coded by ideological bias.
+As shown, the least similar Grokipedia articles exhibit a relative rightward shift in sourcing; however, they remain predominantly left-leaning.
 3.6 Comparison within Topics
-Because these articles cover a diverse set of subject areas, we additionally classified each article using GPT-5 (see Methods for details).
+Because these articles cover a diverse set of subject areas, we additionally classified each article (see Methods for details).
 This allows us to compare similarity and bias shift not only across individual articles but also across broader thematic domains.
@@
 (a) Mean combined similarity scores between Wikipedia and Grokipedia articles within each topic, with standard-error bars.
-(b) Mean political bias shift (Grokipedia – Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia.
+(b) Mean political bias shift (Grokipedia Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia.
 Error bars show standard errors within each topic.
 4 Discussion and Conclusion
-This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 17,700 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles.
+This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across 17,790 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles.
 Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style.
@@
 These omissions substantially reduce navigational affordances and make Grokipedia articles harder to scan, explore, and contextualize.
-The similarity distributions reveal additional nuance.
-The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply.
-This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation.
-In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it departs substantially from the source material, generating content that is stylistically and semantically distinct.
 These results also resonate with earlier linguistic analyses of Wikipedia itself.
@@
 Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
+The similarity distributions reveal additional nuance.
+The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply.
+This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation.
+In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it diverges, generating content that is stylistically and semantically distinct.
 Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [apnews2025, businessinsider2025, yasseri2025conversation].
-Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025].
-Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density.
-In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution.
-Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
+Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia—often copied or paraphrased—while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025].
+Our quantitative results support this view: Grokipedia articles fall into two distinct clusters, with some heavily reliant on Wikipedia pages and others exhibiting systematic divergence.
 As argued by yasseri2025conversation, Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable.
@@
 The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
-Complementing these textual and stylistic findings, our reference-level analysis (Figures 65) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity.
+Complementing these textual and stylistic findings, our reference-level analysis (Figure 6) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity.
 When Grokipedia reproduces Wikipedia closely (high similarity), the resulting entries show little or no bias shift.
-However, low-similarity articles—those most substantially rewritten—display both greater variance and a systematic rightward tendency in bias shift.
-This pattern suggests that editorial divergence in Grokipedia is concentrated precisely where generative rewriting occurs.
-The corresponding breakdown of reference sources further supports this interpretation.
-In the low-similarity subset, citation networks are dominated by a narrower set of news outlets with identifiable political leanings, as measured by the News Media Bias and Factuality dataset [sanchez2024mapping].
-Whereas Wikipedia’s reference base remains diverse and balanced, Grokipedia’s alternative sourcing emphasizes ideologically marked or partisan media domains.
-Together, these results indicate that Grokipedia’s generative process may amplify latent source bias when departing from human-authored baselines, substituting breadth and attribution with ideologically skewed curation.
+However, low-similarity articles display both greater variance and a rightward tendency in bias shift.
 The topic-level analysis reveals that cross-platform divergence is not evenly distributed but concentrated in specific domains.
@@
 This alignment between low similarity and increased ideological skew suggests that Grokipedia’s deviations are not merely stylistic but often substantively directional.
+In effect, Grokipedia trades collective accountability for computational authority.
+Even though some articles do diverge from Wikipedia, they still exhibit a left-leaning bias within (albeit less than their Wikipedia counterparts).
+However, the fact that this rightward shift is more pronounced in specific contested articles might point to the fact that this is not the result of the LLM model generating more neutral sources but a systematic decision by the creators.
+This cannot be examined as the prompts and the parameters used by the model is not publicly available.
+Together, these findings point to a deeper epistemological contrast.
+Wikipedia’s openness renders bias visible and contestable through edits, disputes, and deliberation [yasseri2025conversation].
+Grokipedia replaces this process with opaque, automated authorship, embedding potential bias within model behavior rather than exposing it to scrutiny.
+Despite its stated corrective aim, Grokipedia functions less as an epistemic alternative than as an AI-mediated reconfiguration of Wikipedia, trading collective accountability for computational authority.
 Several limitations qualify these findings.
-First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which around 17,700 have counterparts on Grokipedia.
+First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which 17,790 have counterparts on Grokipedia.
 This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [yasseri2012dynamics].
@@
 The challenge ahead is therefore not only technical, improving factuality or retrieval, but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge.
+As AI-generated reference content may itself be used to train future language models, these patterns have implications beyond a single platform.
+Without transparency, accountability, and sustained human oversight, such derivative systems risk propagating existing biases into the epistemic foundations of future AI systems.
 Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust.
 Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether Grokipedia’s expansion of text corresponds to an expansion, or erosion, of understanding.
-5 Acknowledgement
-This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.
+Acknowledgement
+This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217 and 18/CRT/6049.
+TY acknowledge support from Workday, Inc.
+Data Availability
+Data used in the study are available at 10.5281/zenodo.18433470.